# Kinetic Flat-Histogram Simulations of Non-Equilibrium Stochastic Processes with Continuous and Discontinuous Phase Transitions

### v1

@@ III.4 First Schlögl Model
which are combined with external sources and sinks of particles $B$ and $C$ that keep their concentrations constant in time, where $\rho=N_{A}/N$, $b=N_{B}/N$, $c=N_{C}/N$, and $N$ is the system size ($b$ and $c$ are constants). The dynamics is entirely described by the concentration $\rho$ of $A$ particles.
The concentration of $A$ particles obeys the following time-evolution equation:
$\frac{d\rho}{dt}=-k_{2}\rho^{2}+(k_{1}b-k_{3})\rho+k_{4}c,$ (26)
where we set $k_{1}b=1$, $k_{2}=1$, and $k_{3}=k_{2}\lambda$. The system exhibits a continuous active-absorbing phase transition for $k_{4}=0$ as the control parameter $\lambda$ increases. Eq. (26), with our settings, predicts the threshold $\lambda_{c}=1$. For $k_{4}\neq 0$, there is no transition driven by the control parameter $\lambda$.
Figure 4: (Color Online) Results for the first Schlögl model. The same as in Fig. (3), except that the system is in the active phase for $\lambda<1$ and in the absorbing phase for $\lambda>1$.
Results for the stationary distribution $P(\rho,\lambda)$ obtained using the kinetic flat-histogram algorithm are shown in Fig. (4). We note that the critical $P(\rho,\lambda_{c})$ approaches the absorbing state with a vanishing derivative, as seen for the CP. As in the active-absorbing phase transition of the CP, we obtain a continuous Shannon entropy.
IV Discontinuous Phase Transitions
In this section, we present our findings for well-known bistable non-equilibrium stochastic systems in the mean-field (one-site) regime. Two of the most studied examples are the second Schlögl model [25, 26, 27, 28] and the Ziff-Gulari-Barshad (ZGB) model [29, 30, 32, 33, 34, 35]. These two stochastic processes provide a test to determine whether the kinetic flat-histogram simulations can capture the physics of discontinuous phase transitions, especially weak ones. In all results, we set the system size to $N=10^{3}$.
IV.1 Second Schlögl Model
Parallel to the first Schlögl model, the second Schlögl model [25, 26, 27, 28] was introduced as a prototype of an open chemical reactive system exhibiting discontinuous phase transitions. It also describes chemical reactions in rarefied gases with external particle sources. In the mean-field (one-site) regime, the reaction channels for the second Schlögl model are
[…]
As in the first Schlögl model, the particles $B$ and $C$ can be added or removed to keep their concentrations constant, where $\rho=N_{A}/N$, $b=N_{B}/N$, $c=N_{C}/N$, with $b$ and $c$ constants and $N$ the system size. Once again, the dynamics are entirely described by the concentration $\rho$ of $A$ particles.
The concentration of $A$ particles obeys the following time-evolution equation:
$\frac{d\rho}{dt}=-k_{2}\rho^{3}+k_{1}b\rho^{2}-k_{3}\rho+k_{4}c,$ (28)
where we set $k_{1}b=1$, $k_{2}=1$, $k_{3}=k_{2}\lambda$, and $k_{4}c=\alpha$. Eq. (28), with our settings, predicts a line of discontinuous phase transitions with thresholds $\lambda_{\alpha}$. The critical line ends at a continuous phase transition with critical threshold $\lambda_{c}$ as $\alpha$ increases. When $\alpha=0$, we obtain the exact threshold of a discontinuous active-absorbing phase transition at $\lambda_{0}=1/4$, where the system is active for $\lambda<\lambda_{0}$ and absorbing otherwise. With our settings, the terminus of the critical line is the critical point at $\alpha_{c}=1/27$ and $\lambda_{c}=1/3$ [38].
Figure 5: (Color Online) Results for the second Schlögl model with $\alpha=0.005$. In panel (a), we show the stationary distribution $P(\rho,\lambda)$ for several values of the control parameter $\lambda$. We note that $P(\rho,\lambda)$ presents two well-defined maxima when approaching the threshold. In panel (b), we show $P(\rho,\lambda)$ for two values of $\lambda$ close to the threshold $\lambda_{\alpha}$, one for each of the two phases. Phase coexistence occurs when the two maxima of $P(\rho,\lambda)$ have the same weight. In panel (c), we compare the expected value of the concentration $\rho$ with SSA simulations performed by increasing and decreasing $\lambda$. The SSA curves form a hysteresis cycle, which is a distinctive feature of discontinuous phase transitions. Finally, in panel (d), we show the Shannon entropy $S(\lambda)$, which is clearly discontinuous. The line in panel (d) is only a guide to the eyes.
We show results for the second Schlögl model in Figs. (5), (6), and (7), where we increase $\alpha$ as we approach the terminus of the critical line. For a strong discontinuous phase transition, shown in Fig. (5), all the distinctive features of a discontinuous phase transition are evident. Even the SSA simulations are able to show the hysteresis cycle. Unfortunately, the hysteresis cycle is sensitive to the number of Monte Carlo steps (or the simulation time) discarded to allow the system to reach the stationary distribution. The stationary distribution presents two clear maxima, and the threshold can be estimated by searching for the $P(\rho,\lambda)$ corresponding to phase coexistence, when the two maxima have the same weight. In addition, the Shannon entropy presents a clear jump close to the threshold.
Figure 6: (Color Online) Results for the second Schlögl model with $\alpha=0.025$. The same as in Fig. (5), except that we are unable to see two clear maxima in the stationary distribution $P(\rho,\lambda)$. However, $\ln P(\rho,\lambda)$ is clearly asymmetric, indicating a deviation from the symmetric parabolic shape. This is a sign of a mild discontinuous phase transition. In addition, the transition is too weak to show a hysteresis cycle, and the Shannon entropy is nearly continuous.
We show results for a mild discontinuous phase transition in Fig. (6). It becomes difficult to identify a discontinuous phase transition using standard methods. We discarded a sufficiently long initial simulation time to allow the system to reach the stationary distribution, so no hysteresis cycle is observed. Also, the jump in the Shannon entropy at the threshold begins to vanish. To observe clearer effects, larger system sizes should be used. However, a distinctive feature still remains: $\ln P(\rho,\lambda)$ is clearly asymmetric, allowing us to identify a mild discontinuous phase transition.
Figure 7: (Color Online) Results for the second Schlögl model with $\alpha=0.035$, close to the critical point at $\alpha_{c}=1/27\sim 0.037$ and $\lambda_{c}=1/3\sim 0.333$. The same as in Fig. (5), except that we are unable to see two clear maxima in the stationary distribution $P(\rho,\lambda)$. However, $\ln P(\rho,\lambda)$ is almost symmetric and close to a parabolic shape, which indicates that we are at the onset of the critical point. Therefore, the transition is too weak to show any hysteresis cycle, and the Shannon entropy is practically continuous.
In Fig. (7), we show results close to the critical point. It is difficult to observe any hysteresis cycle because the importance sampling tunnels easily through phase space. In addition, the Shannon entropy is almost continuous. The only evidence of a discontinuous phase transition is the asymmetry of $\ln P(\rho,\lambda)$ near the maxima, indicating a slight deviation from the parabolic shape. If we increase $\alpha$ a little more, $\ln P(\rho,\lambda)$ close to the maxima begin to display a parabolic, symmetric shape.
IV.2 Ziff-Gulari-Barshad Model
Another convenient test of kinetic flat-histogram algorithm is the ZGB model [29, 30, 32, 33, 34, 35] which describes a catalytic metallic surface oxidation by the Langmuir-Hinschelwood process. The Langmuir-Hinschelwood process is the succession of three steps for a chemical reaction: (1) adsorption of the reactants by a catalytic surface, (2) recombination of the reactants, and (3) desorption of the products.
The reaction channels of the ZGB model are:
$C\xrightarrow{k_{1}}A,\text{ }2C\xrightarrow{k_{2}}2B,\text{ }A+B\xrightarrow{k_{3}}2C,\text{ and }A\xrightarrow{k_{4}}C,$ (29)
where $A$, $B$, and $C$ represent a carbon monoxide molecule, an oxygen atom, and an empty surface site, respectively. In the following, we explain each of the channels in Eq. (29):
  • The first channel represents the adsorption of a carbon monoxide molecule, which occupies an empty site;
  • The second channel describes the adsorption of an oxygen molecule, where each of the two resulting atoms occupies an empty site;
  • The third channel corresponds to the oxidation of carbon monoxide, resulting in a carbon dioxide molecule that is ejected from the surface, leaving two empty sites;
  • Finally, the last channel introduces the possibility of a carbon monoxide molecule being ejected from the surface without recombination, leaving a single empty site.
This last channel allows for the existence of a critical line with a terminus at a critical point, as seen in the second Schlögl model.
Let $x_{1}$, $x_{2}$, and $x_{3}$ being the concentrations of particles $A$, $B$, and $C$, respectively. The coupled time-evolution equations for the concentrations are given by
$\left\{\begin{split}&\frac{dx_{1}}{dt}=k_{1}x_{3}-k_{4}x_{1}-k_{3}x_{1}x_{2},\\ &\frac{dx_{2}}{dt}=2k_{2}x_{3}^{2}-k_{3}x_{1}x_{2},\\ &\frac{dx_{3}}{dt}=-k_{1}x_{3}+k_{4}x_{1}-2k_{2}x_{3}^{2}+2k_{3}x_{1}x_{2}.\\ \end{split}\right.$ (30)
@@ IV.2 Ziff-Gulari-Barshad Model
One can see that only two of the three equations are independent. The three concentrations are bounded and in fact, we can describe the dynamics by analyzing the carbon monoxide concentration, where we define $\rho\equiv x_{1}$. Furthermore, we can set $k_{1}=\lambda$, $k_{2}=0.5$, $k_{3}=1$, and $k_{4}=\alpha$. Eq. (30) predicts a line of discontinuous phase transitions with thresholds $\lambda_{\alpha}$ ending in a critical point. With our settings, the terminus of the critical line can be calculated numerically, yielding $\alpha_{c}\sim 0.08$ [38].
Figure 8: (Color Online) Results for the ZGB model with $\alpha=0.02$. In panel (a), we show the stationary distribution $P(\rho,\lambda)$ of the carbon monoxide concentration $\rho$ absorbed by the surface for several values of the control parameter $\lambda$. We note that $P(\rho,\lambda)$ presents two well-defined maxima as the threshold is approached. In panel (b), we show $P(\rho,\lambda)$ for two values of $\lambda$ close to the threshold $\lambda_{\alpha}$, one for each of the two phases. Phase coexistence occurs when the two maxima of $P(\rho,\lambda)$ have the same weight. In panel (c), we compare the expected value of the concentration $\rho$ with SSA simulations performed by increasing and decreasing $\lambda$. The SSA curves form a hysteresis cycle, which is a distinctive feature of discontinuous phase transitions. Finally, in panel (d), we show the Shannon entropy $S(\lambda)$, which is clearly discontinuous. The line in panel (d) is only a guide to the eyes.
@@
We show results for the ZGB model in Figs. (8), (9), and (10). The behavior of the stationary distribution $P(\rho,\lambda)$ is similar to that observed for the second Schlögl model as the critical point is approached. Far from the critical point, as in Fig. (8), we observe a strong discontinuous phase transition, where hysteresis cycles can be easily seen using standard methods. The Shannon entropy displays a clear step, and the threshold can be estimated by identifying the $P(\rho,\lambda)$ with two maxima of equal weight.
Figure 9: (Color Online) Results for the ZGB model with $\alpha=0.06$. The same as in Fig. (8), except that we are almost unable to see two clear maxima in the stationary distribution $P(\rho,\lambda)$. However, $\ln P(\rho,\lambda)$ is clearly asymmetric, indicating a deviation from the symmetric parabolic shape. This corresponds to a mild discontinuous phase transition. In addition, the transition is too weak to show a hysteresis cycle, and the Shannon entropy is nearly continuous.
@@
Eventually, the hysteresis cycle disappears as the critical point is approached by increasing $\alpha$. Also, the step in the Shannon entropy should eventually vanish. However, the asymmetry of $\ln P(\rho,\lambda)$ still persists even at the onset of the critical point, as seen in Fig. (9). In Fig. (10), we show the behavior of $\ln P(\rho,\lambda)$ close to the critical point. We note that $\ln P(\rho,\lambda)$ is symmetric and very close to a parabolic shape.
@@
Figure 10: (Color Online) Results for the ZGB model with $\alpha_{c}=0.08$ (critical point). The same as in Fig. (8), except that $\ln P(\rho,\lambda)$ is nearly symmetric and very close to a parabolic shape. Therefore, there is no hysteresis cycle, and the Shannon entropy is practically continuous.
@@ V Conclusions
We introduced the kinetic flat-histogram algorithm, which is effective for identifying weak discontinuous phase transitions in ergodic stochastic processes. The algorithm performs a random walk in phase space to estimate the stationary distribution of an observable in a non-equilibrium stochastic process.
The kinetic flat-histogram algorithm estimates the stationary distribution of an observable in a stochastic process by rejection sampling, generating trial moves from standard Monte Carlo techniques with an acceptance probability inversely related to the estimated stationary distribution.
The stationary distribution is refined during the random walk through a modification factor, which introduces non-Markovian dynamics. The process becomes Markovian as the estimated stationary distribution converges to the true one. A visitation histogram is accumulated, and the modification factor is updated when this histogram satisfies a flatness condition. Simulation ends when the modification factor reaches a threshold value close to unity.
@@ VI Acknowledgments
We would like to thank CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior), CNPq (Conselho Nacional de Desenvolvimento Científico e tecnológico), and FAPEPI (Fundação de Amparo à Pesquisa do Estado do Piauí) for the financial support. We acknowledge the Dietrich Stauffer Computational Physics Lab., Teresina, Brazil, and Laboratório de Física Teórica e Modelagem Computacional - LFTMC, Teresina, Brazil, where we performed the numerical simulations.

### v2

@@ IV.2 Ziff-Gulari-Barshad Model
Only two of the three equations are independent. The three concentrations are bounded, and the dynamics can be described by analyzing the carbon monoxide concentration, where we define $\rho\equiv x_{1}$. Furthermore, we set $k_{1}=\lambda$, $k_{2}=0.5$, $k_{3}=1$, and $k_{4}=\alpha$. Eq. (34) predicts a line of discontinuous phase transitions with thresholds $\lambda_{\alpha}$ ending at a critical point. With these settings, the terminus of the critical line can be calculated numerically, yielding $\alpha_{c}\sim 0.08$ [71].
@@
We present results for the ZGB model in Figs. (11), (12), and (13). The behavior of the stationary distribution $P(\rho,\lambda)$ is similar to that observed for the second Schlögl model as the critical point is approached. Far from the critical point, as in Fig. (11), we observe a strong discontinuous phase transition, where hysteresis cycles can be easily identified using standard methods. The Shannon entropy displays a clear step, and the threshold can be estimated by identifying the $P(\rho,\lambda)$ with two maxima of equal weight.
Figure 11: (Color Online) Results for the ZGB model with $\alpha=0.02$. Panel (a) shows the stationary distribution $P(\rho,\lambda)$ of the carbon monoxide concentration $\rho$ adsorbed on the surface for several values of the control parameter $\lambda$. We observe that $P(\rho,\lambda)$ exhibits two well-defined maxima as the threshold is approached. Panel (b) presents $P(\rho,\lambda)$ for two values of $\lambda$ close to the threshold $\lambda_{\alpha}$, one for each phase. Phase coexistence occurs when the two maxima of $P(\rho,\lambda)$ have equal weight. Panel (c) compares the expected value of the concentration $\rho$ with SSA simulations performed by increasing and decreasing $\lambda$. The SSA curves form a hysteresis cycle, which is a distinctive feature of discontinuous phase transitions. Finally, panel (d) shows the Shannon entropy $S(\lambda)$, which is clearly discontinuous. The line in panel (d) is shown only as a guide to the eye.
@@
As the critical point is approached by increasing $\alpha$, the hysteresis cycle eventually disappears. The step in the Shannon entropy also vanishes. Nevertheless, the asymmetry of $\ln P(\rho,\lambda)$ persists even near the critical point, as seen in Fig. (12). In Fig. (13), we show the behavior of $\ln P(\rho,\lambda)$ close to the critical point, where it becomes symmetric and nearly parabolic.
Figure 12: (Color Online) Results for the ZGB model with $\alpha=0.06$. As in Fig. (11), it becomes difficult to observe two distinct maxima in the stationary distribution $P(\rho,\lambda)$. However, $\ln P(\rho,\lambda)$ is clearly asymmetric, indicating a deviation from the symmetric parabolic shape. This provides evidence of a weak discontinuous phase transition, where the entropy jump is barely detectable. Additionally, the transition is too weak to exhibit a hysteresis cycle, and the Shannon entropy is practically continuous.
@@
Figure 13: (Color Online) Results for the ZGB model at the critical point $\alpha_{c}=0.08$. As in Fig. (11), $\ln P(\rho,\lambda)$ is now nearly symmetric and very close to a parabolic shape. Consequently, there is no hysteresis cycle, and the Shannon entropy is practically continuous.
@@ V Conclusions
We introduced the kinetic flat-histogram algorithm, which is effective for identifying weak discontinuous phase transitions in ergodic stochastic processes. The algorithm performs a random walk in phase space to estimate the stationary distribution of an observable in a non-equilibrium stochastic process. The kinetic flat-histogram algorithm estimates the stationary distribution of an observable by rejection sampling, generating trial moves using standard Monte Carlo techniques with an acceptance probability inversely related to the estimated stationary distribution.
The stationary distribution is refined during the random walk through a modification factor, which introduces non-Markovian dynamics. The process becomes Markovian as the estimated stationary distribution converges to its asymptotic form. A visitation histogram is accumulated, and the modification factor is updated when this histogram satisfies a flatness condition. The simulation ends when the modification factor reaches a threshold value close to unity.
@@
Similar to the Wang-Landau algorithm, the kinetic flat-histogram algorithm is also prone to error saturation. However, this issue can be mitigated by employing strategies such as the $1/t$ method or adaptive refinement schemes. A theoretical analysis providing a definitive proof of convergence when using $1/t$ updates is still lacking and remains an open question for future research.
@@ VI Acknowledgments
We gratefully acknowledge financial support from CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior), CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico), and FAPEPI (Fundação de Amparo à Pesquisa do Estado do Piauí). We thank the Dietrich Stauffer Computational Physics Lab, Teresina, Brazil, and the Laboratório de Física Teórica e Modelagem Computacional (LFTMC), Teresina, Brazil, for providing computational resources for the numerical simulations. R. S. Ferreira acknowledges support from FAPEMIG (grant FAPEMIG-APQ-06611-24).